OpenStudy (anonymous):

HELP PLZZZZZZZZZZZZZZZ two doves of the same mass are flying towards each other at 5m/s and 7m/s. After they collide, at what speed and in what direction do they move?

OpenStudy (anonymous):

If they are the same size and stick together, they will go at 2 m/s in the direction originally traveled by the duck going 7 m/s. Odd question.

OpenStudy (anonymous):

Apply conservation of momentum to mathematically show that's a bit off! ^_^  \[m_1=m_2=m\ \\ v_{1f}=v_{2f}=v_f\] \[ p_i = p_f\] \[m_1v_{1i}+m_2v_{2i}=m_1v_{1f}+m_2v_{2f}\] \[m(7m/s)+m(-5m/s)=mv_f+mv_f\] \[7m/s-5m/s = 2v_f\] \[2m/s=2v_f\] \[v_f=1m/s\ \ \text{in the positive x direction (the direction of the faster dove)} \] So they will each have a velocity of 1m/s ^_^